A wheel of radius 14.0 cm starts from rest and turns through 2.0 revolutions




in 3.0s.




a. What is its average velocity?




b. What is the tangential velocity of a point on the rim of the wheel?

1
Expert's answer
2021-12-09T15:48:16-0500

Given:

"R=14.0\:\rm cm"

"\omega_0=0"

"N=2"

"t=3.0\:\rm s"


a. The average angular velocity

"\omega=\frac{2\pi N}{t}=\frac{4\pi}{3.0}=4.2\:\rm rad/s"

b. The tangential velocity

"v=\omega R=4.2*14.0=59\:\rm cm/s"